An object of mass 126.309172 g executes SHM with an amplitude of 12.426183 cm and a period 14.856335 s. Find the force acting on the object at time 1.487176 s.
Answer #1

As force = mass * acceleration

Acceleration in SHM, a = -2 A sin (t) (here we are measuring the time from the moment, particle was at the point where acceleration was 0)

So force = mass * ( -2 A sin (t))

= (0.126309172) * ( -((2*3.14)/(14.856335))2 (0.12426183) sin (((2*3.14)/(14.856335))*1.487176)) (here = 2/T)

= -0.00164997 N